OpenStudy (anonymous):

Is there a way to find the third side of a triangle when the other two sides are given?

OpenStudy (lgbasallote):

is it a right triangle?

OpenStudy (anonymous):

The question is: If two sides of a triangle are 6 cm and 8 cm then length of the third side is a. 7 cm b. 2cm c. greater than 2 cm and less than 14 cm d. none of these.

OpenStudy (anonymous):

If 10 is the third side, then it would be a right triangle.

OpenStudy (anonymous):

So is it none of these?

OpenStudy (unklerhaukus):

yes there is The cosine forumla \[c^2=a^2+b^2-2ab\cos C\] in a right angled triangle C (angle opposite hypotenuse c is 90° , cos90°=0

OpenStudy (lgbasallote):

a + b > c where a and b are two legs and c is the third side....you can find a range i guess

OpenStudy (anonymous):

No

OpenStudy (anonymous):

But trig is not involved.

OpenStudy (accessdenied):

If you only know two sides, you can only find a range of values for the side using Triangle Side Inequality theorem.

OpenStudy (lgbasallote):

....which @lgbasallote stated :p

OpenStudy (asnaseer):

 imagine the 6cm side pivotted, and then rotate it all the way to see what the possible lengths of the third side could be.

OpenStudy (unklerhaukus):